# Optimization in Theory and Practice thanks: The full version of the paper can be accessed at https://arxiv.org/abs/0000.00000

### v1

@@ Abstract
Mathematics is used in optimization both to study the properties of problem formulations and to devise algorithms for solving these problems and analyze the properties of these algorithms. These strands are intertwined. The study of formulations includes the existence and characterization of solutions; the stability of solutions under perturbations of the problem data; and duality, which provides an alternative perspective on a given formulation that can shed light on the problem and can sometimes be exploited by algorithms. Algorithms for optimization problems are typically iterative, generating a sequence of candidate solutions that (ideally) converges to a solution for the problem at hand. Algorithms are rigorous in the sense that they allow certain claims about their properties to be proved, but within the rigorous framework, they contain heuristics to guide certain decisions, for example, the choice of pivot in a simplex algorithm for linear programming from among a (potentially enormous) list of valid choices. Along with mathematical expertise and imagination, a good deal of intuition and domain knowledge for the key application areas is required for successful algorithm design, adaptation, and application. Development of software to implement these algorithms requires many other areas of expertise to be brought to bear, involving broader mathematical knowledge and software engineering expertise. The mathematical issues include the vagaries of finite-precision arithmetic, robustness in the face of ill-posed formulations, the interface with numerical linear algebra, and parallel implementation, among many others. Engineering issues (some of which can also be analyzed mathematically) include management of hierarchical memory and data movement.
@@
The focus of this paper is the analysis of algorithms for two classes of continuous optimization problems — those in which the variables are vectors or matrices involving real numbers that are allowed to vary continuously, not being restricted to certain discrete values such as the integers. We consider the convergence and complexity properties of these methods. Are the iterates guaranteed to converge to points with the characteristics of a solution? Can we bound the amount of computation needed to find an approximate solution over all problems in a class — and how do we measure “computation,” given that the cost of obtaining information about functions can vary widely from one application to the next? Can we lower-bound the amount of computation needed for algorithms of a certain type on the hardest problems in a certain class?
@@
We focus especially on the interplay between the theoretical properties of algorithms and their practical performance, which has driven a great deal of research in optimization over the past several decades. For some problem classes, the practical performance of methods on typical problems in the class mirrors the theory well. In other classes, there is a wide gap between theory and practice, with practical performance being much faster on typical problems than the theory predicts. Such gaps can sometimes be narrowed by refining the theory, perhaps by narrowing the problem class in an interesting way. An recent development of this type has occurred in the minimization of smooth nonconvex functions (see Subsection˜5.6). In its full generality, this problem is intractable, but many subclasses of nonconvex problems have been identified (particularly in machine learning) for which global minima can be found efficiently. In other cases of wide gaps between theory and practice, further analysis and computation can improve our understanding of the nature of the gap. For example, we may ask: Are the loose bounds provided by the theory due to rare worst-case instances? Can we quantify the rarity of these instances? Can we mollify these instances, for example by showing that a nearby instance is easy to solve? Gaps may prompt a search for new kinds of algorithms with better theoretical and possibly even better practical properties. A striking instance of this type occurred in linear programming (see Section˜4), where the search for algorithms with polynomial complexity led eventually to the development of interior-point methods, which have proved to be important in practice for solving large-scale problems.
@@
Other notions of approximate optimality including the optimality gap $f(x)-f^{*}$, where $f^{*}$ is the value of $f$ at the solution, and $\mbox{\rm dist}(x,{\cal S})$, where $\cal S$} is the set of solutions to Eq.˜2.1. Neither condition is “checkable" in general, as we need to know the solutions before evaluating them. Nevertheless they are useful in the sense that it may be possible to obtain a bound on the computational effort required to satisfy these conditions approximately to within some positive tolerance $\epsilon$, that is, $f(x)-f^{*}\leq\epsilon$ and $\mbox{\rm dist}(x,{\cal S})\leq\epsilon$, respectively.
@@
We can define approximate solutions of the linear programs Eq.˜2.2 and Eq.˜2.4 in terms of approximate satisfaction of the optimality conditions Eq.˜2.3. One particular measure to which we refer in Section˜4 asssumes that the feasibility conditions (the first three conditions in Eq.˜2.3 are satisfied exactly, while the quantity $\mu$ defined by
@@
In this paper, we focus mostly non-asymptotic analysis, in which we “globalize” the local analysis and seek to say something about the rate of convergence of the algorithm from its initial point. Although it has a long history, this kind of analysis has become much more prominent in recent years. Non-asymptotic analysis of a particular algorithm identifies some scalar quantity $\tau_{k}$ associated with each iteration or each oracle (see below), such that $\tau_{k}\to 0$ as we approach a point satisfying desired optimality conditions. For the unconstrained problem Eq.˜2.1, $\tau_{k}$ might be $f(x^{k})-f^{*}$, $\|\nabla f(x^{k})\|$, or $\mbox{\rm dist}(x^{k},{\cal S})$, where $x^{k}$ is the $k$th iterate of the algorithm. Depending on the algorithm, convergence of $\tau_{k}$ to zero can occur at arithmetic, “sublinear” rates, such as $\tau_{k}\leq A/k$, $\tau_{k}\leq A/\sqrt{k}$, or $\tau_{k}\leq A/k^{2}$ for some constant $A>0$, which depends on the problem and often on the starting point. It can occur at “linear” (sometimes called “exponential”) rates such as $\tau_{k}\leq A(1-\phi)^{k}$, for some constants $A>0$ and $\phi\in(0,1)$. It could occur at “superlinear” rates, for which $\tau_{k+1}/\tau_{k}\to 0$ as $k\to\infty$.
@@
For nonlinear problems such as Eq.˜2.1, the oracle complexity model of Nemirovski and Yudin [73] is widely used to bound the amount of computation required by a certain algorithm on a given class of problems. For a given algorithm, this method defines some “unit of information" (oracle) that the algorithm requires about the function $f$, with many queries made during the course of the algorithm. For one algorithm (e.g. a gradient descent method), a suitable oracle might by the value of $(f(x),\nabla f(x))$ at a given point $x$. For another method (e.g. a stochastic gradient method), the appropriate oracle might be an unbiased estimate of $\nabla f(x)$ with a certain known variance, for a given $x$. The “oracle complexity” of a method is a bound on the number of oracles required to obtain an $\epsilon$-accurate solution. When each iteration of an algorithm requires one oracle (or a fixed number of oracles), there is an obvious relationship between oracle complexity and iteration complexity. This relationship is not always so tight; some algorithms (for example, a gradient-descent method with backtracking) might require a variable number of oracles per iteration.
@@
  • The algorithm design and analysis is based on assumptions about the problem that are much too conservative for most instances, or are tight on only a small fraction of the search space navigated by the algorithm. For instance, an assumption that the gradient of $f$ in Eq.˜2.1 is Lipschitz continuous with some constant $L$ is common in gradient-based methods for this problem. But the local geometry of the problem might see the algorithm making much faster progress in some regions of the parameter space, where the problem behaves better than the worst-case assumptions would suggest.
@@
  • The problems of interest in a given class actually belong to a subclass with properties that set them apart from the wider class and make them easier to solve. One example is the “benign nonconvexity” phenomenon, which has been encountered in many problems (especially from machine learning) over the past 10 years, where global minima of nonconvex objectives are usually found easily [96], despite global minimization of general nonconvex objectives being intractable.
@@
Classically, under certain conditions on the choice of stepsizes, and subject to existence of a solution satisfying (2.3), the running average of the iterates generated by (4.11) converge to an $\epsilon$-accurate solution in $O(1/\epsilon)$ iterations. Since LP has a property called sharpness (essentially that the a certain measure of primal-dual optimality grows linearly with distance to the solution set), this rate can be improved to linear, that is, $O(|\log\epsilon|)$ iterations for an $\epsilon$-accurate solution. However, the constant multiplier of $|\log\epsilon|$ is problem-dependent and almost always extremely large — it is related to the Hoffman constant $H(A,b,c)$ of the system of linear equalities and inequalities that defines the solution, namely,
@@
Convex quadratics serve as a kind of sanity check on gradient-based methods for unconstrained minimization. Such methods correspond to iterative methods for solving the square symmetric linear system Hx=gHx=-g (corresponding to $\nabla f(x)=0$) and can be analyzed with the tools of linear algebra. However, algorithmic strategies such as multistep descent and conjugate gradients developed for convex quadratics often need to be modified significantly for case of more general linear functions, and their convergence properties may carry over only in a weaker form, or not at all.
@@
is known as the $k$th Lanczos subspace for $H$ generated by $r^{0}$. In general, the dimension of this space grows by $1$ at each iteration, until a minimizer is found — in at most $n$ iterations. For search directions $p^{k}$, CG sets $x^{k+1}=x^{k}+\alpha_{k}p^{k}$ for $\alpha_{k}$ that exactly solves $\alpha_{k}=\arg\min_{\alpha>0}\,f(x^{k}+\alpha p^{k})$, The set of directions $\{p^{0},p^{1},\dotsc\}$ is constructed so that $p^{0}=r^{0}$, ${\cal K}_{k}(H;r^{0})=\mbox{span}\{p^{0},p^{1},\dotsc,p^{k}\}$, and the property of conjugacy holds with respect to $H$, that is, $(p^{i})^{T}Hp^{j}=0$ when $i\neq j$. Each $p^{k}$ is generated at iteration $k$, according to the formula
@@
We discuss below alternative approaches to proving convergence and to designing new algorihtms, that have been developed in the past 15 years.
@@
The fundamental idea of momentum that underlies the accelerated gradient methods described above has been enormously influential, though the the great number of specific algorithms based on this idea are rarely implemented as written. Momentum translates into each step being a weighted average of all gradients encountered so far (or since the latest restart), so the philosophy underlying momemtum and gradient averaging is the same.
@@
where ${\cal B}_{k}\subset\{1,2,\dotsc,N\}$ is a “minibatch” with $|{\cal B}_{k}|\ll N$, chosen at random, so that Eq.˜5.24 is an unbiased estimate of $\nabla f(x^{k})$. Classical stochastic-gradient (“SGD”) methods [88] take steps along directions Eq.˜5.24 at each iteration, with steplengths chosen carefully sequences that guarantee convergence in expectation (see for example [72] or [109, Chapter 5]). The currently dominant variant of SGD is Adam [59]. Adam adds two devices to the basic SGD scheme: weighted averaging over gradients encountered at previous iterations, and elementwise scaling of the components of the search direction, the scaling also being learned over all previous iterations. Given the enormous reach of Adam in the training of neural networks, it can be said that the idea of momentum in gradient methods is having a great impact on the practice of optimization in machine learning applications.
@@
The iteration complexity results of these Newton-based algorithms can be converted to oracle complexities. If we define the “oracle” to include explicit evaluation of $\nabla^{2}f(x)$, as well as the gradient $\nabla f(x)$ and function value $f(x)$, the transition is trivial. A more challenging task is to devise oracle complexities that rely only on evaluation of first-order derivatives of $f$. The key observation here is that the Hessian-vector product $\nabla^{2}f(x)p$ for any $x,p\in\mathbb{R}^{n}$ be be obtained by differentiating the inner product $\nabla f(x)^{T}p$ with respect to $x$. If reverse-mode automatic differentiation is used, the cost is bounded by five times the cost of evaluating the gradient $\nabla f(x)$ itself; see for example [43]. Thus, a kind of oracle complexity can be derived by counting the number of gradient evaluations plus the number of Hessian-vector products, and adding the cost of evaluating the function $f$ (usually cheaper). The algorithms in [89, 24] require $\tilde{O}(\epsilon^{-7/4})$ such operations in total. An important element of the proof of this result is to show that no more than $O(\epsilon^{-1/2})$ iterations of conjugate gradient are required within each iteration of the modified Newton method in order to compute a direction that leads to significant progress in reducing $f$.
@@
Although they use only first-order information, quasi-Newton methods typically require much more storage ($O(n^{2})$, to store the matrices $H_{k}$ or $B_{k}$) than the gradient, conjugate gradient, and accelerated gradient methods discussed above, which require storage of just a few vectors of length $n$. The limited-memory BFGS method of [66] (known universally as “L-BFGS”) reduces these storage requirements dramatically by maintaining update vectors $(s^{k},y^{k})$ from only the past $m$ iterations, where $m\ll n$. (Values $m=5$ or $m=10$ are typical.) At iteration $j+1$, initializing with $H_{j+1}^{0}=\left[(y^{j})^{T}s^{j}/(y^{j})^{T}y^{j}\right]\,I$, the update formulae Eq.˜5.29 is applied for $k=j-m+1,j-m+2,\dotsc,j$ to obtain the approximation $H_{j+1}$, with the new search direcion being set to $s^{j+1}=-H_{j+1}\nabla f(x^{j+1})$. Importantly, the matrix $H_{j+1}$ does not need to be stored explicitly; $s^{j+1}$ can be calculated efficiently in $O(mn)$ operations from the vectors $(s^{k},y^{k})$ for $k=j-m+1,\dotsc,j$. Theory for L-BFGS is limited. A safeguarded version can be shown to have similar asymptotic convergence properties as gradient descent, and superlinear local convergence is not seen in general. However, L-BFGS is widely used and effective in practice, often outperforming nonlinear conjugate gradient.

### v2

@@ Abstract
Mathematics is used in optimization to study the properties of problem formulations and to devise algorithms and analyze their properties. These strands are intertwined. The study of formulations includes the existence and characterization of solutions; the stability of solutions under perturbations of the problem data; and duality, which provides an alternative perspective on a given formulation that can shed light on the problem and can sometimes be exploited by algorithms. Algorithms for optimization problems are typically iterative, generating a sequence of candidate solutions that (ideally) converges to a solution for the problem at hand. Algorithms are rigorous in the sense that they allow certain claims about their properties to be proved, but within the rigorous framework, they contain heuristics to guide certain decisions, for example, the choice of pivot in a simplex algorithm for linear programming from among a (potentially enormous) list of valid choices. Along with mathematical expertise and imagination, a good deal of intuition and domain knowledge for the key application areas is required for successful algorithm design, adaptation, and application. Development of software to implement these algorithms requires many other areas of expertise to be brought to bear, involving broader mathematical knowledge and software engineering expertise. The mathematical issues include the vagaries of finite-precision arithmetic, robustness in the face of ill-posed formulations, the interface with numerical linear algebra, and parallel implementation, among many others. Engineering issues (some of which can also be analyzed mathematically) include management of hierarchical memory and data movement.
@@
The focus of this paper is the analysis of algorithms for two classes of continuous optimization problems — problems in which the variables are vectors or matrices involving real numbers that are allowed to vary continuously, not being restricted to certain discrete values such as the integers. We consider the convergence and complexity properties of these methods. Are the iterates guaranteed to converge to points with the characteristics of a solution? Can we bound the amount of computation needed to find an approximate solution over all problems in a class — and how do we measure “computation,” given that the cost of obtaining information about functions can vary widely from one application to the next? Can we lower-bound the amount of computation needed for algorithms of a certain type on the hardest problems in a certain class?
@@
We focus especially on the interplay between the theoretical properties of algorithms and their practical performance, which has driven a great deal of research in optimization over the past several decades. For some problem classes, the practical performance of methods on typical problems in the class mirrors the theory well. In other classes, there is a wide gap between theory and practice, with practical performance being much faster on typical problems than the theory predicts. Such gaps can sometimes be narrowed by refining the theory, perhaps by narrowing the problem class in an interesting way. A recent development of this type has occurred in the minimization of smooth nonconvex functions (see Subsection˜5.6). In its full generality, this problem is intractable, but many subclasses of nonconvex problems have been identified (particularly in machine learning) for which global minima can be found efficiently. In other cases of wide gaps between theory and practice, further analysis and computation can improve our understanding of the nature of the gap. For example, we may ask: Are the loose bounds provided by the theory due to rare worst-case instances? Can we quantify the rarity of these instances? Can we mollify these instances, for example by showing that a nearby instance is usually easy to solve? Gaps may prompt a search for new kinds of algorithms with better theoretical and possibly even better practical properties. A striking instance of this type has occurred in linear programming (see Section˜4), where the search for algorithms with polynomial complexity led eventually to the development of interior-point methods, which have proved to be important in practice for solving large-scale problems.
@@
Other notions of approximate optimality include the optimality gap $f(x)-f^{*}$, where $f^{*}$ is the value of $f$ at the solution, and $\mbox{\rm dist}(x,{\cal S})$, where $\cal S$} is the set of solutions to Eq.˜2.1. Neither condition is “checkable" in general, as we need to know the solutions before evaluating them. Nevertheless they are useful in the sense that it may be possible to obtain a bound on the computational effort required to satisfy these conditions approximately to within some positive tolerance $\epsilon$, that is, $f(x)-f^{*}\leq\epsilon$ and $\mbox{\rm dist}(x,{\cal S})\leq\epsilon$, respectively.
@@
We can define approximate solutions of the linear programs Eq.˜2.2 and Eq.˜2.4 in terms of approximate satisfaction of the optimality conditions Eq.˜2.3. One particular measure to which we refer in Section˜4 asssumes that the feasibility conditions (the first three conditions in Eq.˜2.3) are satisfied exactly, while the quantity $\mu$ defined by
@@
In this paper, we focus mostly on non-asymptotic analysis, in which we “globalize” the local analysis and seek to say something about the rate of convergence of the algorithm from its initial point. Although it has a long history, this kind of analysis has become much more prominent in recent years. Non-asymptotic analysis of a particular algorithm identifies some scalar quantity $\tau_{k}$ associated with each iteration or each oracle (see below), such that $\tau_{k}\to 0$ as we approach a point satisfying desired optimality conditions. For the unconstrained problem Eq.˜2.1, $\tau_{k}$ might be $f(x^{k})-f^{*}$, $\|\nabla f(x^{k})\|$, or $\mbox{\rm dist}(x^{k},{\cal S})$, where $x^{k}$ is the $k$th iterate of the algorithm. Depending on the algorithm, convergence of $\tau_{k}$ to zero can occur at arithmetic, “sublinear” rates, such as $\tau_{k}\leq A/k$, $\tau_{k}\leq A/\sqrt{k}$, or $\tau_{k}\leq A/k^{2}$ for some constant $A>0$, which depends on the problem and often on the starting point. It can occur at “linear” (sometimes called “exponential” or “geometric") rates such as $\tau_{k}\leq A(1-\phi)^{k}$, for some constants $A>0$ and $\phi\in(0,1)$. It could occur at “superlinear” rates, for which $\tau_{k+1}/\tau_{k}\to 0$ as $k\to\infty$.
@@
For nonlinear problems such as Eq.˜2.1, the oracle complexity model of Nemirovski and Yudin [73] is widely used to bound the amount of computation required by a certain algorithm on a given class of problems. For a given algorithm, this method defines some “unit of information" (oracle) that the algorithm requires about the function $f$, with many queries made during the course of the algorithm. For one algorithm (e.g. a gradient descent method), a suitable oracle might by the value of $(f(x),\nabla f(x))$ at a given point $x$. For another method (e.g. a stochastic gradient method), the appropriate oracle might be an unbiased random estimate of $\nabla f(x)$ with a certain known variance, for a given $x$. The “oracle complexity” of a method is a bound on the number of oracles required to obtain an $\epsilon$-accurate solution. When each iteration of an algorithm requires one oracle (or a fixed number of oracles), there is an obvious relationship between oracle complexity and iteration complexity. This relationship is not always so tight; some algorithms (for example, a gradient-descent method with backtracking) might require a variable number of oracles per iteration.
@@
  • The algorithm design and analysis is based on assumptions about the problem that are much too conservative for most instances, or are tight on only a small fraction of the search space navigated by the algorithm. For instance, an assumption that the gradient of $f$ in Eq.˜2.1 is Lipschitz continuous with some constant $L$ is common in gradient-based methods for this problem. But the local geometry of the problem might see the algorithm making much faster progress in some regions of the parameter space, where the problem has much nicer properties than the worst-case assumptions would suggest.
@@
  • The problems of interest in a given class actually belong to a subclass with properties that set them apart from the wider class and make them easier to solve. One example is the “benign nonconvexity” phenomenon, which has been encountered in many problems (especially from machine learning) over the past 10 years, where global minima of certain nonconvex objectives are usually found easily [96], despite global minimization of general nonconvex objectives being intractable.
@@
Classically, under certain conditions on the choice of stepsizes, and subject to existence of a solution satisfying (2.3), the running average of the iterates generated by (4.11) converges to an $\epsilon$-accurate solution in $O(1/\epsilon)$ iterations. Since LP has a property called sharpness (essentially that a certain measure of primal-dual optimality grows linearly with distance to the solution set), this rate can be improved to linear, that is, $O(|\log\epsilon|)$ iterations for an $\epsilon$-accurate solution. However, the constant multiplier of $|\log\epsilon|$ is problem-dependent and almost always extremely large — it is related to the Hoffman constant $H(A,b,c)$ of the system of linear equalities and inequalities that defines the solution, namely,
@@
Convex quadratics serve as a kind of sanity check on gradient-based methods for unconstrained minimization. Such methods correspond to iterative methods for solving the square symmetric linear system Hx=gHx=-g (corresponding to $\nabla f(x)=0$) and can be analyzed with the tools of linear algebra. However, algorithmic strategies such as multistep descent and conjugate gradients developed for convex quadratics often need to be modified significantly for case of more general nonlinear functions, and their convergence properties may carry over only in a weaker form, or not at all.
@@
is known as the $k$th Krylov subspace for $H$ generated by $r^{0}$. In general, the dimension of this space grows by $1$ at each iteration, until a minimizer is found — in at most $n$ iterations. For search directions $p^{k}$, CG sets $x^{k+1}=x^{k}+\alpha_{k}p^{k}$ for $\alpha_{k}$ that exactly solves $\alpha_{k}=\arg\min_{\alpha>0}\,f(x^{k}+\alpha p^{k})$, The set of directions $\{p^{0},p^{1},\dotsc\}$ is constructed so that $p^{0}=r^{0}$, ${\cal K}_{k}(H;r^{0})=\mbox{span}\{p^{0},p^{1},\dotsc,p^{k}\}$, and the property of conjugacy holds with respect to $H$, that is, $(p^{i})^{T}Hp^{j}=0$ when $i\neq j$. Each $p^{k}$ is generated at iteration $k$, according to the formula
@@
We discuss below alternative approaches to proving convergence and to designing new algorithms, that have been developed in the past 15 years.
@@
The fundamental idea of momentum that underlies the accelerated gradient methods described above has been enormously influential, though the great number of specific algorithms based on this idea are rarely implemented as written. Momentum translates into each step being a weighted average of all gradients encountered so far (or since the latest restart), so the philosophy underlying momemtum and gradient averaging is the same.
@@
where ${\cal B}_{k}\subset\{1,2,\dotsc,N\}$ is a “minibatch” with $|{\cal B}_{k}|\ll N$, chosen at random, so that Eq.˜5.24 is an unbiased estimate of $\nabla f(x^{k})$. Classical stochastic-gradient (“SGD”) methods [88] take steps along directions Eq.˜5.24 at each iteration, with steplengths chosen to guarantee convergence in expectation (see for example [72] or [109, Chapter 5]). The currently dominant variant of SGD is Adam [59]. Adam adds two devices to the basic SGD scheme: weighted averaging over gradients encountered at previous iterations, and elementwise scaling of the components of the search direction, with the scaling also being learned over all previous iterations. Given the enormous reach of Adam in the training of neural networks, it can be said that the idea of momentum in gradient methods is having a great impact on the practice of optimization in machine learning applications.
@@
The iteration complexity results of these Newton-based algorithms can be converted to oracle complexities. If we define the “oracle” to include explicit evaluation of $\nabla^{2}f(x)$, as well as the gradient $\nabla f(x)$ and function value $f(x)$, the transition is trivial. A more challenging task is to devise oracle complexities that rely only on evaluation of first-order derivatives of $f$. The key observation here is that the Hessian-vector product $\nabla^{2}f(x)p$ for any $x,p\in\mathbb{R}^{n}$ can be obtained by differentiating the inner product $\nabla f(x)^{T}p$ with respect to $x$. If reverse-mode automatic differentiation is used, the cost is bounded by five times the cost of evaluating the gradient $\nabla f(x)$ itself; see for example [43]. Thus, a kind of oracle complexity can be derived by counting the number of gradient evaluations plus the number of Hessian-vector products, and adding the cost of evaluating the function $f$ (usually cheaper). The algorithms in [89, 24] require $\tilde{O}(\epsilon^{-7/4})$ such operations in total. An important element of the proof of this result is to show that no more than $O(\epsilon^{-1/2})$ iterations of conjugate gradient are required within each iteration of the modified Newton method in order to compute a direction that leads to significant progress in reducing $f$.
@@
Although they use only first-order information, quasi-Newton methods typically require much more storage ($O(n^{2})$, to store the matrices $H_{k}$ or $B_{k}$) than the gradient, conjugate gradient, and accelerated gradient methods discussed above, which require storage of just a few vectors of length $n$. The limited-memory BFGS method of [66] (known universally as “L-BFGS”) reduces these storage requirements dramatically by maintaining update vectors $(s^{k},y^{k})$ from only the past $m$ iterations, where $m\ll n$. (Values $m=5$ or $m=10$ are typical.) At iteration $j+1$, initializing with $H_{j+1}^{0}=\left[(y^{j})^{T}s^{j}/(y^{j})^{T}y^{j}\right]\,I$, the update formulae Eq.˜5.29 is applied for $k=j-m+1,j-m+2,\dotsc,j$ to obtain the approximation $H_{j+1}$, with the new search direction being set to $s^{j+1}=-H_{j+1}\nabla f(x^{j+1})$. Importantly, the matrix $H_{j+1}$ does not need to be stored explicitly; $s^{j+1}$ can be calculated efficiently in $O(mn)$ operations from the vectors $(s^{k},y^{k})$ for $k=j-m+1,\dotsc,j$. Theory for L-BFGS is limited. A safeguarded version can be shown to have similar asymptotic convergence properties as gradient descent, and superlinear local convergence is not seen in general. However, L-BFGS is widely used and effective in practice, often outperforming nonlinear conjugate gradient.